# Designing quantum chemistry algorithms with just-in-time compilation

### v4

@@ 4.1. Compilation Overhead
| (ss—ss)/(11—11) | 1Q1T | 284.8 | 25 | ||||
| (ss—ss)/(33—33) | 1Q1T | 512.0 | 74 | ||||
| (pp—pp)/(11—11) | 1Q1T | 638.8 | 75 | ||||
| (pp—pp)/(33—33) | 1Q1T | 880.6 | 94 | ||||
| (dd—dd)/(11—11) | 1QnT | 6011 | 194 | ||||
| (ff—ff)/(11—11) | 1QnT | 7697 | 270 | ||||
| (gg—gg)/(11—11) | 1QnT | 59468 | 602 |
4.2. Benchmark with GPU4PySCF and TeraChem
In this work, integrals for the Coulomb (J) and exchange (K) terms are screened simultaneously. GPU4PySCF and xQC use the same screening technique. The number of significant shell quartets in each JK calculation is identical in both cases. In double-precision, xQC reproduces GPU4PySCF’s results exactly, so we omit a detailed FP64 accuracy analysis. Under our single precision scheme, JK matrix elements are computed in single precision but accumulated in FP64; all other operations including diagonalization and direct inversion of the iterative subspace (DIIS)—remain in FP64.
We use a customized TeraChem v1.9 build with f-orbital support. Although both GPU4PySCF and TeraChem employ the Schwarz inequality to bound integral magnitudes, they apply it at different levels: TeraChem at the primitive shell level and GPU4PySCF at the contracted level. Their SCF initial guesses and DIIS settings also differ, so our GPU4PySCF–TeraChem comparisons are not strictly apples-to-apples. For all calculations shown, the screening threshold for both J and K matrices is set to . SCF with GPU4PySCF/xQC converges in 14 iterations, whereas the same SCF with TeraChem converges in 11 iterations. This does not imply TeraChem’s settings are superior, as convergence behavior can vary with the system. A sample TeraChem input is provided in Appendix A. All the basis sets are in Cartesian format.
When both codes run in FP64, their total energies agree to within 0.001 mHa. When all integrals are evaluated in FP32, we observe uncertainties on the order of 1 mHa in both xQC and TeraChem (Table 4), reinforcing the well-known result that pure single-precision arithmetic is insufficient for high-accuracy quantum-chemical calculations. A more practical mixed-precision algorithm is among our future directions.
@@ 4.2. Benchmark with GPU4PySCF and TeraChem
As shown in Table 5, xQC delivers a twofold speedup over GPU4PySCF v1.4 when using the 6-31G* basis set and a fourfold gain with the larger def2-TZVPP basis. Although TeraChem v1.9 generally outpaces GPU4PySCF v1.4, its advantage is only marginal. In double-precision, xQC outperforms TeraChem by about 4x when def2-TZVPP basis set is used, and a less significant 1.6x for 6-31G* basis. In single-precision, xQC slightly outperforms TeraChem on the small basis set and runs twice as fast on def2-TZVPP. Both TeraChem and xQC still employ double precision for eigenvalue decomposition and DIIS, making these linear-algebra routines a significant cost on the A10-24 GB GPU. Isolating the JK kernel, xQC is approximately three times faster than TeraChem.
| Molecule | Protocol | GPU4PySCF | TeraChem | xQC | |||
@@
| gly30 | HF/6-31G*/A100-80G/FP64 | 1.9 s | 28.1 s | 1.3 s | 19.08 s | 0.8 s | 12.6 s |
| gly30 | HF/def2-TZVPP/A100-80G/FP64 | 78.4 s | 1125.7 s | 91.4 s | 1055.1 s | 18.8 s | 268.1 s |
| gly30 | HF/6-31G*/A10-24G/FP32 | - | - | 1.0 s | 23.12 s | 0.7 s | 18.08 s |
| gly30 | HF/def2-TZVPP/A10-24G/FP32 | - | - | 63.5 s | 870.81 s | 21.2 s | 423.19 s |
5. Conclusion
@@
6. Future Work and Limitations
High level abstractions of JIT. Although our current implementation relies heavily on handwritten CUDA kernels, it would be preferable to express these routines in a high-level language such as Triton (TKC, 19), Numba (LPS, 15), or JAX (BFH+, 18). These frameworks offer richer support for kernel fusion, making it easier to embed quantum-chemistry operations within machine-learning pipelines. For example, one could fuse the on-the-fly computation of integrals with subsequent matrix–matrix multiplications (not only contraction with density matrices), yielding more efficient end-to-end workflows. Further, one could write a customized kernel for the derivatives with respect to contraction coefficients and exponents, then port to an existing auto-differentiation framework (ZC, 22).
Supporting more functionalities. A practical quantum‐chemistry workflow requires more than the acceleration of Coulomb and exchange kernels. The exchange-correlation functional in DFT, the gradients and Hessians of Coulomb, exchange and exchange-correlation terms, post Hartree-Fock methods, and a wide array of molecular properties are all essential. By integrating xQC with GPU4PySCF, users can build sophisticated simulation pipelines, and any of these routines can be further accelerated on demand using JIT compilation.
Mixed precision algorithms. This paper focuses on the performance of the double precision and single precision separately. Thanks to the Schwartz inequality, the computational load of the integrals can be naturally split into FP64 calculations and FP32 calculations without loss of accuracy theoretically. Other research efforts (LUM, 11; MM, 13) show that 90% of the computations can be in FP32. A naive mixed precision strategy can be easily proposed based on this work. One should notice, even though the value of integral can be represented in FP32, not every underlying arithmetic operation for computing the integral can necessarily be computed with FP32. More effort is needed to rigorously implement a strategy without error accumulation.
Screening techniques. The intrinsic scalings of the Coulomb and exchange terms are different. The number of significant Coulomb terms increases quadratically, while the number of significant exchange terms increases linearly with respect to the system size. It is more efficient to evaluate the Coulomb and exchange terms separately. In some quantum chemistry packages such as TeraChem, the Coulomb terms are efficiently evaluated with J-engine (WH, 96). Besides the Schwartz inequality, which gives a rigorous upper bound, other estimates can further eliminate the unnecessary calculations. For example, the MBIE (LDO, 05), QQR (MLFO, 12), and CSAM (TO, 17). MBIE provides a rigorous upperbound, but it is expensive to calcuate. QQR and CSAM take advantage of distance dependence in their estimates, hence improving the performance for large systems. But those estimates are less rigorous.
@@ 6. Future Work and Limitations
Xiaojie Wu led the conceptualization of JIT algorithms and developed xQC v0.1, and wrote the initial draft. Qiming Sun contributed the improvements of the JK kernels v1.0 to v1.4 in GPU4PySCF, which are not documented elsewhere. Yuanheng Wang explored single-precision algorithms in GPU4PySCF, discussed JIT algorithms in xQC v0.1, and wrote the manuscript.
Appendix A TeraChem input sample
In this paper, we use the following sample TeraChem script. The coordinates, basis, and precision arguments are changed correspondingly.
run energy method hf basis 6-31gs precision single coordinates gly30.xyz charge 0 timings yes threcl 1e-13 threex 1e-13 guess sad xtol 1e-6 purify no fock incremental
Appendix B Analyze the bottlenecks of 1q1t algorithm and 1qnt algorithm
We benchmark the 1q1t algorithm on the gly120 system with the following customized basis set in NWChem format (PAD+, 19),
X S
0.2700058226E+00 1
0.2700058226E+00 1
0.2700058226E+00 1
0.2700058226E+00 1
0.2700058226E+00 1
0.2700058226E+00 1
The same basis set is assigned to all types of elements, for profiling purpose only. The number of primitives varies in the tests. The 1q1t algorithm is largely memory bound when the number of primitives is small. For a (ss—ss) kernel with uncontracted orbitals, when switching from FP64 to FP32, the speedup of FP64/FP32 is only 2. The computing intensity increases rapidly with the number of primitives. When the numbers of primitives of all the shells are 4, the FP32 to FP64 speedup achieves the theoretical peak performance of A10-24G, which has 32:1 FP64 / FP32 compute units. However, when the number of primitives is greater than 5, we observed the register spill out onto local memory. This is essentially due to the compiler (NVRTC/NVCC) not unrolling the nested loops of primitives. There are multiple ways to manually unroll the loops, but for simplicity of the code, we leave it for future work. The similar issue also occurs in the case of other patterns, such as , , and so on.
The performance bottleneck in the 1qnt algorithm arises from two distinct phases: integral generation and contraction. In the integral generation phase, the cost of calculating the integral values is independent of the number of density matrices. In the contraction phase, the integral value is contracted with every density matrix, so its cost grows linearly with the matrix count. To quantify this behavior, we vary the number of density matrices and recorded the total wall-clock time (Figure 7). We constructed the following basis set in NWChem format for the test, which assigns the same d and f orbtials to all types of elements:
X D
1.2700058226E+00 1
X F
1.2700058226E+00 1
The results show an almost perfectly linear relation between the overall cost and the number of density matrices. For the (dd—dd) pattern, integral generation requires 96 ms and each contraction 67 ms; in a higher-angular-momentum case, these times rise to 1,518 ms and 837 ms, respectively. Consequently, when only one density matrix is used (e.g., in restricted HF/DFT methods), the bottleneck lies in integral generation, whereas with two or more matrices (for example, unrestricted HF/DFT methods), density matrix contraction dominates. In the latter scenario, more sophisticated contraction strategies are needed, such as tensor core-accelerated kernels.
We emphasize that this analysis is approximate: increased register pressure from handling multiple density matrices can further affect kernel performance, and those effects are not reflected in our current timing estimates.
Acknowledgements.
@@
The authors appreciate the support of Dr. Wen Yan and the ByteDance Seed management team. Special thanks are also extended to Dr. Zhichen Pu for his discussions. The authors also thank the PySCF community for providing valuable tools that contributed to this work, especially the GPU4PySCF contributors.References
- AAB+ [15] Martín Abadi, Ashish Agarwal, Paul Barham, Eugene Brevdo, Zhifeng Chen, Craig Citro, Greg S. Corrado, Andy Davis, Jeffrey Dean, Matthieu Devin, Sanjay Ghemawat, Ian Goodfellow, Andrew Harp, Geoffrey Irving, Michael Isard, Yangqing Jia, Rafal Jozefowicz, Lukasz Kaiser, Manjunath Kudlur, Josh Levenberg, Dandelion Mané, Rajat Monga, Sherry Moore, Derek Murray, Chris Olah, Mike Schuster, Jonathon Shlens, Benoit Steiner, Ilya Sutskever, Kunal Talwar, Paul Tucker, Vincent Vanhoucke, Vijay Vasudevan, Fernanda Viégas, Oriol Vinyals, Pete Warden, Martin Wattenberg, Martin Wicke, Yuan Yu, and Xiaoqiang Zheng. TensorFlow: Large-scale machine learning on heterogeneous systems, 2015. Software available from tensorflow.org.
- AAF+ [10] Andrey Asadchev, Veerendra Allada, Jacob Felder, Brett M. Bode, Mark S. Gordon, and Theresa L. Windus. Uncontracted rys quadrature implementation of up to g functions on graphical processing units. Journal of Chemical Theory and Computation, 6(3):696–704, 2010. PMID: 26613300.
- AV [24] Andrey Asadchev and Edward F. Valeev. 3-center and 4-center 2-particle gaussian ao integrals on modern accelerated processors. The Journal of Chemical Physics, 160(24):244109, 06 2024.
- Bec [14] Axel D. Becke. Perspective: Fifty years of density-functional theory in chemical physics. The Journal of Chemical Physics, 140(18):18A301, 04 2014.
- BFH+ [18] James Bradbury, Roy Frostig, Peter Hawkins, Matthew James Johnson, Chris Leary, Dougal Maclaurin, George Necula, Adam Paszke, Jake VanderPlas, Skye Wanderman-Milne, and Qiao Zhang. JAX: composable transformations of Python+NumPy programs, 2018.
- DP [75] James D. Dill and John A. Pople. Self-consistent molecular orbital methods. xv. extended gaussian-type basis sets for lithium, beryllium, and boron. J. Chem. Phys., 62:2921–2923, 1975.
- DRK [76] Michel Dupuis, John Rys, and Harry F. King. Evaluation of molecular integrals over gaussian basis functions. The Journal of Chemical Physics, 65(1):111–116, 07 1976.
- EGF+ [21] Evgeny Epifanovsky, Andrew T. B. Gilbert, Xintian Feng, Joonho Lee, Yuezhi Mao, Narbe Mardirossian, Pavel Pokhilko, Alec F. White, Marc P. Coons, Adrian L. Dempwolff, Zhengting Gan, Diptarka Hait, Paul R. Horn, Leif D. Jacobson, Ilya Kaliman, Jörg Kussmann, Adrian W. Lange, Ka Un Lao, Daniel S. Levine, Jie Liu, Simon C. McKenzie, Adrian F. Morrison, Kaushik D. Nanda, Felix Plasser, Dirk R. Rehn, Marta L. Vidal, Zhi-Qiang You, Ying Zhu, Bushra Alam, Benjamin J. Albrecht, Abdulrahman Aldossary, Ethan Alguire, Josefine H. Andersen, Vishikh Athavale, Dennis Barton, Khadiza Begam, Andrew Behn, Nicole Bellonzi, Yves A. Bernard, Eric J. Berquist, Hugh G. A. Burton, Abel Carreras, Kevin Carter-Fenk, Romit Chakraborty, Alan D. Chien, Kristina D. Closser, Vale Cofer-Shabica, Saswata Dasgupta, Marc de Wergifosse, Jia Deng, Michael Diedenhofen, Hainam Do, Sebastian Ehlert, Po-Tung Fang, Shervin Fatehi, Qingguo Feng, Triet Friedhoff, James Gayvert, Qinghui Ge, Gergely Gidofalvi, Matthew Goldey, Joe Gomes, Cristina E. González-Espinoza, Sahil Gulania, Anastasia O. Gunina, Magnus W. D. Hanson-Heine, Phillip H. P. Harbach, Andreas Hauser, Michael F. Herbst, Mario Hernández Vera, Manuel Hodecker, Zachary C. Holden, Shannon Houck, Xunkun Huang, Kerwin Hui, Bang C. Huynh, Maxim Ivanov, Ádám Jász, Hyunjun Ji, Hanjie Jiang, Benjamin Kaduk, Sven Kähler, Kirill Khistyaev, Jaehoon Kim, Gergely Kis, Phil Klunzinger, Zsuzsanna Koczor-Benda, Joong Hoon Koh, Dimitri Kosenkov, Laura Koulias, Tim Kowalczyk, Caroline M. Krauter, Karl Kue, Alexander Kunitsa, Thomas Kus, István Ladjánszki, Arie Landau, Keith V. Lawler, Daniel Lefrancois, Susi Lehtola, Run R. Li, Yi-Pei Li, Jiashu Liang, Marcus Liebenthal, Hung-Hsuan Lin, You-Sheng Lin, Fenglai Liu, Kuan-Yu Liu, Matthias Loipersberger, Arne Luenser, Aaditya Manjanath, Prashant Manohar, Erum Mansoor, Sam F. Manzer, Shan-Ping Mao, Aleksandr V. Marenich, Thomas Markovich, Stephen Mason, Simon A. Maurer, Peter F. McLaughlin, Maximilian F. S. J. Menger, Jan-Michael Mewes, Stefanie A. Mewes, Pierpaolo Morgante, J. Wayne Mullinax, Katherine J. Oosterbaan, Garrette Paran, Alexander C. Paul, Suranjan K. Paul, Fabijan Pavošević, Zheng Pei, Stefan Prager, Emil I. Proynov, Ádám Rák, Eloy Ramos-Cordoba, Bhaskar Rana, Alan E. Rask, Adam Rettig, Ryan M. Richard, Fazle Rob, Elliot Rossomme, Tarek Scheele, Maximilian Scheurer, Matthias Schneider, Nickolai Sergueev, Shaama M. Sharada, Wojciech Skomorowski, David W. Small, Christopher J. Stein, Yu-Chuan Su, Eric J. Sundstrom, Zhen Tao, Jonathan Thirman, Gábor J. Tornai, Takashi Tsuchimochi, Norm M. Tubman, Srimukh Prasad Veccham, Oleg Vydrov, Jan Wenzel, Jon Witte, Atsushi Yamada, Kun Yao, Sina Yeganeh, Shane R. Yost, Alexander Zech, Igor Ying Zhang, Xing Zhang, Yu Zhang, Dmitry Zuev, Alán Aspuru-Guzik, Alexis T. Bell, Nicholas A. Besley, Ksenia B. Bravaya, Bernard R. Brooks, David Casanova, Jeng-Da Chai, Sonia Coriani, Christopher J. Cramer, György Cserey, III DePrince, A. Eugene, Jr. DiStasio, Robert A., Andreas Dreuw, Barry D. Dunietz, Thomas R. Furlani, III Goddard, William A., Sharon Hammes-Schiffer, Teresa Head-Gordon, Warren J. Hehre, Chao-Ping Hsu, Thomas-C. Jagau, Yousung Jung, Andreas Klamt, Jing Kong, Daniel S. Lambrecht, WanZhen Liang, Nicholas J. Mayhall, C. William McCurdy, Jeffrey B. Neaton, Christian Ochsenfeld, John A. Parkhill, Roberto Peverati, Vitaly A. Rassolov, Yihan Shao, Lyudmila V. Slipchenko, Tim Stauch, Ryan P. Steele, Joseph E. Subotnik, Alex J. W. Thom, Alexandre Tkatchenko, Donald G. Truhlar, Troy Van Voorhis, Tomasz A. Wesolowski, K. Birgitta Whaley, III Woodcock, H. Lee, Paul M. Zimmerman, Shirin Faraji, Peter M. W. Gill, Martin Head-Gordon, John M. Herbert, and Anna I. Krylov. Software for the frontiers of quantum chemistry: An overview of developments in the q-chem 5 package. The Journal of Chemical Physics, 155(8):084801, 08 2021.
- FL [08] N. Flocke and V. Lotrich. Efficient electronic integrals and their generalized derivatives for object oriented implementations of electronic structure calculations. Journal of Computational Chemistry, 29(16):2722–2736, 2008.
- FTS+ [16] M. J. Frisch, G. W. Trucks, H. B. Schlegel, G. E. Scuseria, M. A. Robb, J. R. Cheeseman, G. Scalmani, V. Barone, G. A. Petersson, H. Nakatsuji, X. Li, M. Caricato, A. V. Marenich, J. Bloino, B. G. Janesko, R. Gomperts, B. Mennucci, H. P. Hratchian, J. V. Ortiz, A. F. Izmaylov, J. L. Sonnenberg, D. Williams-Young, F. Ding, F. Lipparini, F. Egidi, J. Goings, B. Peng, A. Petrone, T. Henderson, D. Ranasinghe, V. G. Zakrzewski, J. Gao, N. Rega, G. Zheng, W. Liang, M. Hada, M. Ehara, K. Toyota, R. Fukuda, J. Hasegawa, M. Ishida, T. Nakajima, Y. Honda, O. Kitao, H. Nakai, T. Vreven, K. Throssell, J. A. Montgomery, Jr., J. E. Peralta, F. Ogliaro, M. J. Bearpark, J. J. Heyd, E. N. Brothers, K. N. Kudin, V. N. Staroverov, T. A. Keith, R. Kobayashi, J. Normand, K. Raghavachari, A. P. Rendell, J. C. Burant, S. S. Iyengar, J. Tomasi, M. Cossi, J. M. Millam, M. Klene, C. Adamo, R. Cammi, J. W. Ochterski, R. L. Martin, K. Morokuma, O. Farkas, J. B. Foresman, and D. J. Fox. Gaussian˜16 Revision C.01, 2016. Gaussian Inc. Wallingford CT.
- HJO [00] Trygve Helgaker, Poul Jørgensen, and Jeppe Olsen. Atomic Basis Functions, chapter 6, pages 201–255. John Wiley & Sons, Ltd, 2000.
- HP [88] Martin Head‐Gordon and John A. Pople. A method for two‐electron gaussian integral and integral derivative evaluation using recurrence relations. The Journal of Chemical Physics, 89(9):5777–5786, 11 1988.
- LA [04] C. Lattner and V. Adve. Llvm: a compilation framework for lifelong program analysis & transformation. In International Symposium on Code Generation and Optimization, 2004. CGO 2004., pages 75–86, 2004.
- LDO [05] Daniel S. Lambrecht, Bernd Doser, and Christian Ochsenfeld. Rigorous integral screening for electron correlation methods. The Journal of Chemical Physics, 123(18):184102, 11 2005.
- LPS [15] Siu Kwan Lam, Antoine Pitrou, and Stanley Seibert. Numba: a llvm-based python jit compiler. In Proceedings of the Second Workshop on the LLVM Compiler Infrastructure in HPC, LLVM ’15, New York, NY, USA, 2015. Association for Computing Machinery.
- LSZC [25] Rui Li, Qiming Sun, Xing Zhang, and Garnet Kin-Lic Chan. Introducing gpu acceleration into the python-based simulations of chemistry framework. The Journal of Physical Chemistry A, 129(5):1459–1468, 2025. PMID: 39846468.
- LUM [11] Nathan Luehr, Ivan S. Ufimtsev, and Todd J. Martínez. Dynamic precision for electron repulsion integral evaluation on graphical processing units (gpus). Journal of Chemical Theory and Computation, 7(4):949–954, 2011. PMID: 26606344.
- MD [78] Larry E McMurchie and Ernest R Davidson. One- and two-electron integrals over cartesian gaussian functions. Journal of Computational Physics, 26(2):218–231, 1978.
- MLFO [12] Simon A. Maurer, Daniel S. Lambrecht, Denis Flaig, and Christian Ochsenfeld. Distance-dependent schwarz-based integral estimates for two-electron integrals: Reliable tightness vs. rigorous upper bounds. The Journal of Chemical Physics, 136(14):144107, 04 2012.
- MM [13] Yipu Miao and Kenneth M. Jr. Merz. Acceleration of electron repulsion integral evaluation on graphics processing units via use of recurrence relations. Journal of Chemical Theory and Computation, 9(2):965–976, 2013. PMID: 26588740.
- Nee [22] F. Neese. Software update: the orca program system, version 5.0. WIRES Comput. Molec. Sci., 12(1):e1606, 2022.
- NVI [25] NVIDIA Corporation. NVRTC: NVIDIA Runtime Compilation Library. NVIDIA Corporation, 2025. Accessed: 2025-06-17.
- OUN+ [17] Ryosuke Okuta, Yuya Unno, Daisuke Nishino, Shohei Hido, and Crissman Loomis. Cupy: A numpy-compatible library for nvidia gpu calculations. In Proceedings of Workshop on Machine Learning Systems (LearningSys) in The Thirty-first Annual Conference on Neural Information Processing Systems (NIPS), 2017.
- PAD+ [19] Benjamin P. Pritchard, Doaa Altarawy, Brett Didier, Tara D. Gibsom, and Theresa L. Windus. A new basis set exchange: An open, up-to-date resource for the molecular sciences community. J. Chem. Inf. Model., 59:4814–4820, 2019.
- PGM+ [19] Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, Alban Desmaison, Andreas Köpf, Edward Yang, Zach DeVito, Martin Raison, Alykhan Tejani, Sasank Chilamkurthy, Benoit Steiner, Lu Fang, Junjie Bai, and Soumith Chintala. PyTorch: an imperative style, high-performance deep learning library. Curran Associates Inc., Red Hook, NY, USA, 2019.
- Sun [15] Qiming Sun. Libcint: An efficient general integral library for gaussian basis functions. Journal of Computational Chemistry, 36(22):1664–1671, 2015.
- TKC [19] Philippe Tillet, H. T. Kung, and David Cox. Triton: an intermediate language and compiler for tiled neural network computations. In Proceedings of the 3rd ACM SIGPLAN International Workshop on Machine Learning and Programming Languages, MAPL 2019, page 10–19, New York, NY, USA, 2019. Association for Computing Machinery.
- TO [17] Travis H. Thompson and Christian Ochsenfeld. Distance-including rigorous upper bounds and tight estimates for two-electron integrals over long- and short-range operators. The Journal of Chemical Physics, 147(14):144101, 10 2017.
- VBa [23] Jorge Luis Galvez Vallejo, Giuseppe M.J. Barca, and Mark S. Gordon and. High-performance gpu-accelerated evaluation of electron repulsion integrals. Molecular Physics, 121(9-10):e2112987, 2023.
- WA [05] Florian Weigend and Reinhart Ahlrichs. Balanced basis sets of split valence, triple zeta valence and quadruple zeta valence quality for h to rn: Design and assessment of accuracy. Phys. Chem. Chem. Phys., 7:3297, 2005.
- WH [96] Christopher A. White and Martin Head‐Gordon. A j matrix engine for density functional theory calculations. The Journal of Chemical Physics, 104(7):2620–2629, 02 1996.
- WHJ+ [24] Yuanheng Wang, Diptarka Hait, K. Grace Johnson, O. Jonathan Fajen, Juncheng Harry Zhang, Rubén D. Guerrero, and Todd J. Martínez. Extending gpu-accelerated gaussian integrals in the terachem software package to f type orbitals: Implementation and applications. The Journal of Chemical Physics, 161(17):174118, 11 2024.
- WSP+ [25] Xiaojie Wu, Qiming Sun, Zhichen Pu, Tianze Zheng, Wenzhi Ma, Wen Yan, Yu Xia, Zhengxiao Wu, Mian Huo, Xiang Li, Weiluo Ren, Sheng Gong, Yumin Zhang, and Weihao Gao. Enhancing gpu-acceleration in the python-based simulations of chemistry frameworks. WIREs Computational Molecular Science, 15(2):e70008, 2025. e70008 CMS-1146.R2.
- ZC [22] Xing Zhang and Garnet Kin-Lic Chan. Differentiable quantum chemistry with pyscf for molecules and materials at the mean-field level and beyond. The Journal of Chemical Physics, 157(20):204801, 11 2022.

### v5

@@ 4.1 Compilation Overhead
| (ss|ss)/(11|11) | 1Q1T | 284.8 | 25 | ||||
| (ss|ss)/(33|33) | 1Q1T | 512.0 | 74 | ||||
| (pp|pp)/(11|11) | 1Q1T | 638.8 | 75 | ||||
| (pp|pp)/(33|33) | 1Q1T | 880.6 | 94 | ||||
| (dd|dd)/(11|11) | 1QnT | 6,011 | 194 | ||||
| (ff|ff)/(11|11) | 1QnT | 7,697 | 270 | ||||
| (gg|gg)/(11|11) | 1QnT | 59,468 | 602 |
4.2 Benchmark with GPU4PySCF and TeraChem
In this work, integrals for the Coulomb (J) and exchange (K) terms are screened simultaneously. GPU4PySCF and JoltQC use the same screening technique. The number of significant shell quartets in each JK calculation is identical in both cases. In double-precision, JoltQC reproduces GPU4PySCF’s results exactly, so we omit a detailed FP64 accuracy analysis. Under our single precision scheme, JK matrix elements are computed in single precision but accumulated in FP64; all other operations including diagonalization and direct inversion of the iterative subspace (DIIS)—remain in FP64.
We use a customized TeraChem v1.9 build with f-orbital support. Although both GPU4PySCF and TeraChem employ the Schwarz inequality to bound integral magnitudes, they apply it at different levels: TeraChem at the primitive shell level and GPU4PySCF at the contracted level. Their SCF initial guesses and DIIS settings also differ, so our GPU4PySCF–TeraChem comparisons are not strictly apples-to-apples. For all calculations shown, the screening threshold for both J and K matrices is set to . SCF with GPU4PySCF/JoltQC converges in 14 iterations, whereas the same SCF with TeraChem converges in 11 iterations. This does not imply TeraChem’s settings are superior, as convergence behavior can vary with the system. A sample TeraChem input is provided in Appendix 7. All the basis sets are in Cartesian format.
When both codes run in FP64, their total energies agree to within 0.001 mHa. When all integrals are evaluated in FP32, we observe uncertainties on the order of 1 mHa in both JoltQC and TeraChem (Table 4), reinforcing the well-known result that pure single-precision arithmetic is insufficient for high-accuracy quantum-chemical calculations. A more practical mixed-precision algorithm is among our future directions.
@@ 4.2 Benchmark with GPU4PySCF and TeraChem
As shown in Table 5, JoltQC delivers a twofold speedup over GPU4PySCF v1.4 when using the 6-31G* basis set and a fourfold gain with the larger def2-TZVPP basis. Although TeraChem v1.9 generally outpaces GPU4PySCF v1.4, its advantage is only marginal. In double-precision, JoltQC outperforms TeraChem by about 4x when def2-TZVPP basis set is used, and a less significant 1.6x for 6-31G* basis. In single-precision, JoltQC slightly outperforms TeraChem on the small basis set and runs twice as fast on def2-TZVPP. Both TeraChem and JoltQC still employ double precision for eigenvalue decomposition and DIIS, making these linear-algebra routines a significant cost on the A10-24 GB GPU. Isolating the JK kernel, JoltQC is approximately three times faster than TeraChem.
| Molecule | Protocol | GPU4PySCF | TeraChem | JoltQC | |||
@@
| gly30 | HF/6-31G*/A100/FP64 | 1.9 s | 28.1 s | 1.3 s | 19.1 s | 0.8 s | 10.4 s |
| gly30 | HF/def2-TZVPP/A100/FP64 | 78.4 s | 1,125.7 s | 91.4 s | 1,055.1 s | 18.8 s | 268.1 s |
| gly30 | HF/6-31G*/A10/FP32 | - | - | 1.0 s | 23.1 s | 0.7 s | 16.5 s |
| gly30 | HF/def2-TZVPP/A10/FP32 | - | - | 63.5 s | 870.8 s | 20.0 s | 406.1 s |
5 Conclusion
@@
6 Future Work and Limitations
High level abstractions of JIT. Although our current implementation relies heavily on handwritten CUDA kernels, it would be preferable to express these routines in a high-level language such as Triton [28], Numba [13], or JAX [5]. These frameworks offer richer support for kernel fusion, making it easier to embed quantum-chemistry operations within machine-learning pipelines. For example, one could fuse the on-the-fly computation of integrals with subsequent matrix–matrix multiplications (not only contraction with density matrices), yielding more efficient end-to-end workflows. Further, one could write a customized kernel for the derivatives with respect to contraction coefficients and exponents, then port to an existing auto-differentiation framework [34].
Supporting more functionalities. A practical quantum‐chemistry workflow requires more than the acceleration of Coulomb and exchange kernels. The exchange-correlation functional in DFT, the gradients and Hessians of Coulomb, exchange and exchange-correlation terms, post Hartree-Fock methods, and a wide array of molecular properties are all essential. By integrating JoltQC with GPU4PySCF, users can build sophisticated simulation pipelines, and any of these routines can be further accelerated on demand using JIT compilation.
Mixed precision algorithms. This paper focuses on the performance of the double precision and single precision separately. Thanks to the Schwartz inequality, the computational load of the integrals can be naturally split into FP64 calculations and FP32 calculations without loss of accuracy theoretically. Other research efforts [17, 20] show that 90% of the computations can be in FP32. A naive mixed precision strategy can be easily proposed based on this work. One should notice, even though the value of integral can be represented in FP32, not every underlying arithmetic operation for computing the integral can necessarily be computed with FP32. More effort is needed to rigorously implement a strategy without error accumulation.
Screening techniques. The intrinsic scalings of the Coulomb and exchange terms are different. The number of significant Coulomb terms increases quadratically, while the number of significant exchange terms increases linearly with respect to the system size. It is more efficient to evaluate the Coulomb and exchange terms separately. In some quantum chemistry packages such as TeraChem, the Coulomb terms are efficiently evaluated with J-engine [32]. Besides the Schwartz inequality, which gives a rigorous upper bound, other estimates can further eliminate the unnecessary calculations. For example, the MBIE [14], QQR [18], and CSAM [27]. MBIE provides a rigorous upperbound, but it is expensive to calcuate. QQR and CSAM take advantage of distance dependence in their estimates, hence improving the performance for large systems. But those estimates are less rigorous.
@@ 6 Future Work and Limitations
Xiaojie Wu led the conceptualization of JIT algorithms and developed JoltQC v0.1, and wrote the initial draft. Qiming Sun contributed the improvements of the JK kernels v1.0 to v1.4 in GPU4PySCF, which are not documented elsewhere. Yuanheng Wang explored single-precision algorithms in GPU4PySCF, discussed JIT algorithms in JoltQC v0.1, and wrote the manuscript.
Acknowledgments
@@
The authors appreciate the support of Dr. Wen Yan and the ByteDance Seed management team. Special thanks are also extended to Dr. Zhichen Pu for his discussions. The authors also thank the PySCF community for providing valuable tools that contributed to this work, especially the GPU4PySCF contributors.